find three points that solve the equation -x-2y=-10

Answer by ikleyn(52781)

.
-x - 2y = - 10.    (1)

Express  

x = 10 - 2y        (2)

from the given equation.


Now, give three arbitrary values for y. For example,

y = 1, 2 3.


Calculate the corresponding values of x using formula (2). You will get

x = 10-2*1 = 8;  x = 10-2*2 = 6  and  x = 10-2*3 = 4.


So, these three points (8,1), (6,2) and (4,3) all lie in the straight line (1).

Solved.